Prove: 12|N iff 4| N ^ 3 | N.


1
Expert's answer
2022-04-13T18:10:32-0400

"Let\\,\\,12|N\\Rightarrow N=12k\\Rightarrow N=3\\left( 4k \\right) ,N=4\\left( 3k \\right) \\Rightarrow 4|N\\land 3|N\\\\Let\\,\\,4|N\\land 3|N\\Rightarrow \\left\\{ \\begin{array}{c}\tN=4k\\\\\tN=3m\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\tN=4k\\\\\t4k=3m\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\tN=4k\\\\\tk=3\\left( m-k \\right)\\\\\\end{array} \\right. \\Rightarrow \\\\\\Rightarrow N=4\\cdot 3\\left( m-k \\right) \\Rightarrow N=12\\left( m-k \\right) \\Rightarrow 12|N"